# Practical Policy Distillation for Reinforcement Learning in Radio Access Networks

### v1

@@ Abstract
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This is the author’s version of the paper accepted for publication in IEEE International Symposium on Personal, Indoor and Mobile Radio Communications, 2025. The final published version is available at IEEE Xplore, https://doi.org/[DOI].
@@ I Introduction
AI is increasingly recognized as a key enabler of improved performance and efficiency in RANs. AI-driven solutions have been explored across a broad range of RAN functions, including channel estimation and symbol decoding at the physical layer [SPM+:19, SoD:22]; fast control loops, such as link adaptation and radio resource management, at the medium access layer [KHV+:22, Chen:23, calabrese:18, SKC:22]; and slower functionalities such as mobility management, load balancing, and network configuration [AoF:23, GSE+:22]. Factors such as user mobility, interference, traffic fluctuations, and the heterogeneity of both RAN infrastructure and user equipment (UE) introduce significant variability. As a result, achieving reliable performance in real-world RAN deployments requires AI models that can adapt to diverse and dynamic network conditions [generalization_2024]. In other words, models must generalize effectively and perform robustly, even on previously unseen data [tu2020understanding].
Training an AI model for a RAN function to generalize across diverse RAN environments—rather than fine-tuning separate models for specific conditions, such as those encountered in a single radio cell—is essential for scalable AI integration in RANs [generalization_2024]. Reducing model proliferation enables more efficient and cost-effective lifecycle management. In contrast, models fine-tuned to specific network conditions often lack the robustness required to adapt to changes in network and environmental dynamics, necessitating frequent updates and compromising performance consistency. However, achieving effective model generalization requires extensive and diverse training data, as well as sufficiently large models capable of learning and extrapolating complex patterns from large datasets [tu2020understanding].
@@
Knowledge distillation [gou2021knowledge] has been proposed to address the challenge of deploying complex models on resource-constrained devices. It enables knowledge transfer from a large, high-performing teacher model to a smaller, computationally efficient student model, with negligible degradation in accuracy [hinton2015distilling]. Originally introduced for supervised learning tasks [BCR:06, hinton2015distilling], recent work [RGG+:15, SoF:19] has extended these techniques to policy distillation in reinforcement learning for control applications. However, the ability of policy distillation to retain the generalization capability of the teacher model—and its applicability to practical problems in RANs—remains relatively unexplored.
@@ III Problem formulation
We consider the problem of LA in the downlink of a RAN, where the transmission rate is dynamically adjusted to match the link capacity by optimizing the modulation and coding scheme (MCS) parameters based on time- and frequency-varying channel conditions and radio interference. In 5G systems, the MCS parameters—comprising the modulation order and code rate—are represented by an MCS index [3rd_generation_partnership_project_3gpp_technical_2024], denoted by integer values $m$ ranging from 0 to M1M-1.
LA algorithms deployed in present-day RANs typically rely on two control loops: an inner loop that maps the latest channel state information (CSI) reported by a UE to an initial signal-to-interference-plus-noise ratio (SINR) estimate, and an outer loop that adjusts this estimate based on hybrid automatic repeat request (HARQ) feedback to maximize spectral efficiency while meeting a target block error rate (BLER). However, the outer loop converges slowly [BGA:16] to a fixed BLER target; therefore, it struggles to adapt to dynamic channel conditions under bursty traffic. In contrast, reinforcement learning (RL)-based approach can learn a policy that predicts the optimal MCS index in real time, offering the potential to overcome the limitations of rule-based LA algorithms.
@@ III-A RL-based Link Adaptation
We model the LA problem as an episodic Markov decision process (MDP) $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right\rangle$, where each episode corresponds to the lifespan of a UE data packet—from its first transmission to either successful reception or packet drop after $N$ transmission attempts, as illustrated in Fig. 1. The action space $\mathcal{A}=\{a\mid a\in\{0,\dots,M-1\}\}$ comprises $M$ discrete values, representing the MCS indices used to adapt the link spectral efficiency. In our evaluations, we consider 28 MCS indices (i.e., $M=28$), as defined in Table 5.1.3.1-2 of the 5G specifications [3rd_generation_partnership_project_3gpp_technical_2024].
@@
Following [DWT+:25], we use a reward function $r_{n}=r(s_{n},a_{n})\in\mathcal{R}$ for each individual packet transmission, defined as
@@
To support model generalization across the RAN environment during training, we consider two key enablers. First, we design the state $s\in\mathcal{S}$ to encapsulate a mixture of semi-static and dynamic information. The semi-static information serves two purposes: (i) to characterize the heterogeneity of the RAN deployment and its configuration relative to the UE—including deployment type, location, orientation, interrelations among network sites or radio cells, and parameters such as antenna array type, carrier frequency, bandwidth, and transmit power—and (ii) to describe the UE’s characteristics (e.g., chipset and capabilities), enabling the model to operate across a diverse set of devices. The dynamic information, in contrast, captures state variables relevant to LA behavior, such as path loss, CSI, HARQ feedback, and many more; see, e.g., [DWT+:25] for the complete list of states.
@@ IV Policy Distillation
Knowledge distillation, referred to as policy distillation in the RL context [RGG+:15], involves transferring a policy learned by an agent (the teacher), through interaction with an environment, into a smaller, simpler model (the student) without significant performance degradation. Policy distillation can be performed in both online and offline settings [RGG+:15, rusu2016policydistillation, SoF:19]. In online policy distillation, the student model learns concurrently with the teacher, as the latter improves its policy through ongoing interaction with the environment. In contrast, offline policy distillation assumes that the teacher has been fully trained. The student is then trained using a dataset generated by the teacher’s fixed policy.
@@ IV-A Offline Policy Distillation
Following [rusu2016policydistillation], we consider a distillation dataset $\mathcal{D}^{T}=\{(s_{i},\mathbf{q}_{i}^{T})\}_{i}$, where each sample $i$ consists of a state $s_{i}\in\mathcal{S}$ from the LA function and a vector of unnormalized Q-values $\mathbf{q}_{i}^{T}=\pi^{T}(s_{i})$ produced by the teacher policy $\pi^{T}$. These Q-values serve as target outputs for training a student policy $\pi^{S}$.
@@
Here, a softmax transformation is applied to both the teacher’s Q-values $\mathbf{q}_{i}^{T}$, scaled by a temperature parameter $\tau>0$, and the student’s Q-values $\mathbf{q}_{i}^{S}$, which are left unscaled. In the context of policy distillation, Rusu et al. [rusu2016policydistillation] recommend using lower values of $\tau$ to sharpen the output distribution. Since Q-values represent the expected future discounted rewards for each possible action, sharpening accentuates the differences in action preferences and improves the fidelity of the transferred policy.
Unlike prior work, our objective is to distill a student policy $\pi^{S}$ that retains the teacher’s generalization capability while being significantly more compact. To this end, we revisit and extend two offline policy distillation strategies: single- and multi-policy distillation [rusu2016policydistillation].

### v2

@@ Abstract
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This is the author’s version of the paper accepted for publication in IEEE International Symposium on Personal, Indoor and Mobile Radio Communications, 2025. The final published version is available at IEEE Xplore, https://doi.org/10.1109/PIMRC62392.2025.11274591.
@@ I Introduction
AI is increasingly recognized as a key enabler of improved performance and efficiency in RANs. AI-driven solutions have been explored across a broad range of RAN functions, including channel estimation and symbol decoding at the physical layer [12, 17]; fast control loops, such as link adaptation and radio resource management, at the medium access layer [11, 6, 5, 15]; and slower functionalities such as mobility management, load balancing, and network configuration [2, 9]. Factors such as user mobility, interference, traffic fluctuations, and the heterogeneity of both RAN infrastructure and user equipment (UE) introduce significant variability. As a result, achieving reliable performance in real-world RAN deployments requires AI models that can adapt to diverse and dynamic network conditions [16]. In other words, models must generalize effectively and perform robustly, even on previously unseen data [19].
Training an AI model for a RAN function to generalize across diverse RAN environments—rather than fine-tuning separate models for specific conditions, such as those encountered in a single radio cell—is essential for scalable AI integration in RANs [16]. Reducing model proliferation enables more efficient and cost-effective lifecycle management. In contrast, models fine-tuned to specific network conditions often lack the robustness required to adapt to changes in network and environmental dynamics, necessitating frequent updates and compromising performance consistency. However, achieving effective model generalization requires extensive and diverse training data, as well as sufficiently large models capable of learning and extrapolating complex patterns from large datasets [19].
@@
Knowledge distillation [8] has been proposed to address the challenge of deploying complex models on resource-constrained devices. It enables knowledge transfer from a large, high-performing teacher model to a smaller, computationally efficient student model, with negligible degradation in accuracy [10]. Originally introduced for supervised learning tasks [4, 10], recent work [13, 18] has extended these techniques to policy distillation in reinforcement learning for control applications. However, the ability of policy distillation to retain the generalization capability of the teacher model—and its applicability to practical problems in RANs—remains relatively unexplored.
@@ III Problem formulation
We consider the problem of LA in the downlink of a RAN, where the transmission rate is dynamically adjusted to match the link capacity by optimizing the modulation and coding scheme (MCS) parameters based on time- and frequency-varying channel conditions and radio interference. In 5G systems, the MCS parameters—comprising the modulation order and code rate—are represented by an MCS index [1], denoted by integer values $m$ ranging from 0 to M1M-1.
LA algorithms deployed in present-day RANs typically rely on two control loops: an inner loop that maps the latest channel state information (CSI) reported by a UE to an initial signal-to-interference-plus-noise ratio (SINR) estimate, and an outer loop that adjusts this estimate based on hybrid automatic repeat request (HARQ) feedback to maximize spectral efficiency while meeting a target block error rate (BLER). However, the outer loop converges slowly [3] to a fixed BLER target; therefore, it struggles to adapt to dynamic channel conditions under bursty traffic. In contrast, reinforcement learning (RL)-based approach can learn a policy that predicts the optimal MCS index in real time, offering the potential to overcome the limitations of rule-based LA algorithms.
@@ III-A RL-based Link Adaptation
We model the LA problem as an episodic Markov decision process (MDP) $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right\rangle$, where each episode corresponds to the lifespan of a UE data packet—from its first transmission to either successful reception or packet drop after $N$ transmission attempts, as illustrated in Fig. 1. The action space $\mathcal{A}=\{a\mid a\in\{0,\dots,M-1\}\}$ comprises $M$ discrete values, representing the MCS indices used to adapt the link spectral efficiency. In our evaluations, we consider 28 MCS indices (i.e., $M=28$), as defined in Table 5.1.3.1-2 of the 5G specifications [1].
@@
Following [7], we use a reward function $r_{n}=r(s_{n},a_{n})\in\mathcal{R}$ for each individual packet transmission, defined as
@@
To support model generalization across the RAN environment during training, we consider two key enablers. First, we design the state $s\in\mathcal{S}$ to encapsulate a mixture of semi-static and dynamic information. The semi-static information serves two purposes: (i) to characterize the heterogeneity of the RAN deployment and its configuration relative to the UE—including deployment type, location, orientation, interrelations among network sites or radio cells, and parameters such as antenna array type, carrier frequency, bandwidth, and transmit power—and (ii) to describe the UE’s characteristics (e.g., chipset and capabilities), enabling the model to operate across a diverse set of devices. The dynamic information, in contrast, captures state variables relevant to LA behavior, such as path loss, CSI, HARQ feedback, and many more; see, e.g., [7] for the complete list of states.
@@ IV Policy Distillation
Knowledge distillation, referred to as policy distillation in the RL context [13], involves transferring a policy learned by an agent (the teacher), through interaction with an environment, into a smaller, simpler model (the student) without significant performance degradation. Policy distillation can be performed in both online and offline settings [13, 14, 18]. In online policy distillation, the student model learns concurrently with the teacher, as the latter improves its policy through ongoing interaction with the environment. In contrast, offline policy distillation assumes that the teacher has been fully trained. The student is then trained using a dataset generated by the teacher’s fixed policy.
@@ IV-A Offline Policy Distillation
Following [14], we consider a distillation dataset $\mathcal{D}^{T}=\{(s_{i},\mathbf{q}_{i}^{T})\}_{i}$, where each sample $i$ consists of a state $s_{i}\in\mathcal{S}$ from the LA function and a vector of unnormalized Q-values $\mathbf{q}_{i}^{T}=\pi^{T}(s_{i})$ produced by the teacher policy $\pi^{T}$. These Q-values serve as target outputs for training a student policy $\pi^{S}$.
@@
Here, a softmax transformation is applied to both the teacher’s Q-values $\mathbf{q}_{i}^{T}$, scaled by a temperature parameter $\tau>0$, and the student’s Q-values $\mathbf{q}_{i}^{S}$, which are left unscaled. In the context of policy distillation, Rusu et al. [14] recommend using lower values of $\tau$ to sharpen the output distribution. Since Q-values represent the expected future discounted rewards for each possible action, sharpening accentuates the differences in action preferences and improves the fidelity of the transferred policy.
Unlike prior work, our objective is to distill a student policy $\pi^{S}$ that retains the teacher’s generalization capability while being significantly more compact. To this end, we revisit and extend two offline policy distillation strategies: single- and multi-policy distillation [14].